OpenStudy (anonymous):

Help, please: (3k^2+1k-2)/(4k-2) divided by (k^2+3k+2)/(2k^2+5k-3)

OpenStudy (mathstudent55):

Is this the problem? \( \dfrac{3k^2 + 1k - 2}{4k - 2} \div \dfrac{k^2 + 3k +2}{2k^2 + 5k - 3} \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

The first step is to turn the division into a multiplication by flipping the second fraction.

OpenStudy (mathstudent55):

\(=\dfrac{3k^2 + 1k - 2}{4k - 2} \times \dfrac{2k^2 + 5k - 3}{k^2 + 3k +2} \)

OpenStudy (mathstudent55):

The next step is to factor both numerators and both denominators.

OpenStudy (anonymous):